# Tension-induced giant actuation in unstructured elastic sheets

### v1

@@ D.2 Benthem’s Laplace transform method
It is only a matter of some (tedious) algebra to calculate the residues in (A26) (using  (A24) and the corresponding definitions of the functions $f_{\pm}(y,p)$ (A25)). Finally, we obtain the result in (A8), with
$C_{k}=$ $-\frac{\sin p_{k}}{\cos^{2}p_{k}}\sum_{m}\frac{(p_{k}\alpha_{m}+\beta_{m})}{\left[p_{k}^{2}-\left(\frac{m\pi}{2}\right)^{2}\right]^{2}}m\pi\sin\left(\frac{m\pi}{2}\right),$ (A27a)
$\hat{C}_{k}=$ $-[f^{(-)}_{p,y}(1,-p_{k})-\sin^{2}p_{k}\,f_{p}^{(-)}(1,-p_{k})]\frac{\sin p_{k}}{\cos^{2}p_{k}},$ (A27b)
@@
$m=2k-1$ for $k\in{\mathbb{Z}}^{+}$; where, to write things more compactly, we have introduced the function $f^{(-)}_{p}(y,p)$
$f^{(-)}_{p}(y,p)=f_{p}(y,p)+e^{-pW/2}\sum_{m}\frac{(p\,\alpha_{m}+\beta_{m})}{\left[p^{2}-\left(\frac{m\pi}{2}\right)^{2}\right]^{2}}\cos\left(\frac{m\pi}{2}y\right).$ (A28)
D.3 Near-threshold analysis
@@
Based on the planar stress field results discussed previously, we examine the buckling of a clamped square sheet, of side $\sim W$ and thickness $t$, subject to a uniform bi-axial stress field of modulus $\sim T$, which is tensile along the $y$ axis, $\sigma_{yy}\sim T$, and compressing the $x$ axis, $\sigma_{xx}\sim-T$.
@@
We assume then an undulating displacement field $\zeta(x,y)=\delta\cdot g_{\lambda}(x,y)$. The amplitude $\delta$ is infinitesimal and $g_{\lambda}(x,y)$ is some undulating function along the compressive axis with a characteristic length $\lambda$, yet unknown. Along the tensile axis, $g_{\lambda}(x,y)$ simply suppresses the amplitude at the edges ±$\pm W/2$. The excess bending energy associated with this perturbation is $U_{\mathrm{bend}}\sim B\int~{}(\nabla^{2}\zeta)^{2}\mathrm{d}A\sim B\,W^{2}\left[\left(\frac{\delta}{\lambda^{2}}\right)^{2}+\left(\frac{\delta}{W^{2}}\right)^{2}\right]$ — the full sheet width at the edges is clamped, and hence going out of plane effectively means that the sheet develops a non-zero curvature along both the tensile and the compressing directions. For a given $\lambda$, the excess stretching energy of this double curvature is $U_{\mathrm{stretch}}\sim\int\sigma:\epsilon~{}\mathrm{d}A\sim TW^{2}\delta^{2}/W^{2}-\sigma_{xx}(\lambda)W^{2}\delta^{2}/\lambda^{2}$, and hence the planar state becomes unstable only if the compressive load exceeds
$\sigma_{xx}(\lambda)\sim\frac{B}{\lambda^{2}}+\left(\frac{B}{W^{4}}+\frac{T}{W^{2}}\right)\lambda^{2}.$ (A29)
However, since $\sigma_{xx}$ is also proportional to $T$, we find that
$T\sim\frac{B}{\lambda^{2}}\frac{\left(1+\frac{\lambda^{4}}{W^{4}}\right)}{\left(1+\frac{\lambda^{2}}{W^{2}}\right)}.$ (A30)
@@
Since the wavelength is limited by $W$, we express $\lambda$ as $\lambda=\xi W$, with $\xi\in(0,1]$; then we can write $T\sim\frac{B}{W^{2}}g(\xi)$, with $g(\xi)=\frac{1}{\xi^{2}}\frac{(1+\xi^{4})}{(1+\xi^{2})}$. For $\xi\ll 1$, $g(\xi)\sim\xi^{-2}>1$; while for $\xi\sim 1$ (i.e. $\xi=1-\varepsilon$, $\varepsilon\ll 1$), $g(\xi)\sim(1-\varepsilon)^{-2}>1$. Thus, we have that $g(\xi<1)>1$ and $g(\xi=1)=1$; or, in other words, $g(\xi=1)=1$ is the minimum value of $g(\xi)$ in the interval $\xi\in(0,1]$. This minimum gives the threshold value for the tension
$T_{c}\sim\frac{B}{W^{2}}.$ (A31)
@@
In terms of the longitudinal in-plane end-stretching imposed, $\epsilon_{\text{planar}}$, we conclude that the end-stretched sheet buckles when the longitudinal in-plane displacement $\epsilon_{\text{planar}}$ exceeds $\epsilon_{c}$, with
$\epsilon_{c}\sim\frac{B}{YW^{2}}.$ (A32)

### v2

@@ Appendix E Details of Mathematical Modelling
It is only a matter of some (tedious) algebra to calculate the residues in (S26) (using  (S24) and the corresponding definitions of the functions $f_{\pm}(y,p)$ (S25)). Finally, we obtain the result in (S8), with
$C_{k}=$ $-\frac{\sin p_{k}}{\cos^{2}p_{k}}\sum_{m}\frac{(p_{k}\alpha_{m}+\beta_{m})}{\left[p_{k}^{2}-\left(\frac{m\pi}{2}\right)^{2}\right]^{2}}m\pi\sin\left(\frac{m\pi}{2}\right),$ (S27a)
$\hat{C}_{k}=$ $-[f^{(-)}_{p,y}(1,-p_{k})-\sin^{2}p_{k}\,f_{p}^{(-)}(1,-p_{k})]\frac{\sin p_{k}}{\cos^{2}p_{k}},$ (S27b)
@@
$m=2k-1$ for $k\in{\mathbb{Z}}^{+}$; here, to write things more compactly, we have introduced the function $f^{(-)}_{p}(y,p)$
$f^{(-)}_{p}(y,p)=f_{p}(y,p)+e^{-pW/2}\sum_{m}\frac{(p\,\alpha_{m}+\beta_{m})}{\left[p^{2}-\left(\frac{m\pi}{2}\right)^{2}\right]^{2}}\cos\left(\frac{m\pi}{2}y\right).$ (S28)
Near-threshold analysis
@@
Based on the planar stress field results discussed previously, we examine the buckling of a clamped square sheet, of side $\sim W$ and thickness $t$, subject to a uniform bi-axial stress field of modulus $\sim T_{0}$, which is tensile along the $y$ axis, $\sigma_{yy}\sim T_{0}$, and compressing the $x$ axis, $\sigma_{xx}\sim-T_{0}$.
@@
We assume then an undulating displacement field $\zeta(x,y)=\delta\cdot g_{\lambda}(x,y)$. The amplitude $\delta$ is infinitesimal and $g_{\lambda}(x,y)$ is some undulating function along the compressive axis with a characteristic length $\lambda$, yet unknown. Along the tensile axis, $g_{\lambda}(x,y)$ simply suppresses the amplitude at the edges ±$\pm W/2$. The excess bending energy associated with this perturbation is $U_{\mathrm{bend}}\sim B\int~{}(\nabla^{2}\zeta)^{2}\mathrm{d}A\sim B\,W^{2}\left[\left(\frac{\delta}{\lambda^{2}}\right)^{2}+\left(\frac{\delta}{W^{2}}\right)^{2}\right]$ — the full sheet width at the edges is clamped, and hence going out of plane effectively means that the sheet develops a non-zero curvature along both the tensile and the compressing directions. For a given $\lambda$, the excess stretching energy of this double curvature is $U_{\mathrm{stretch}}\sim\int\sigma:\epsilon~{}\mathrm{d}A\sim T_{0}W^{2}\delta^{2}/W^{2}-\sigma_{xx}(\lambda)W^{2}\delta^{2}/\lambda^{2}$, and hence the planar state becomes unstable only if the compressive load exceeds
$\sigma_{xx}(\lambda)\sim\frac{B}{\lambda^{2}}+\left(\frac{B}{W^{4}}+\frac{T_{0}}{W^{2}}\right)\lambda^{2}.$ (S29)
However, since $\sigma_{xx}$ is also proportional to $T_{0}$, we find that
$T_{0}\sim\frac{B}{\lambda^{2}}\frac{\left(1+\frac{\lambda^{4}}{W^{4}}\right)}{\left(1+\frac{\lambda^{2}}{W^{2}}\right)}.$ (S30)
@@
Since the wavelength is limited by $W$, we express $\lambda$ as $\lambda=\xi W$, with $\xi\in(0,1]$; then we can write $T_{0}\sim\frac{B}{W^{2}}g(\xi)$, with $g(\xi)=\frac{1}{\xi^{2}}\frac{(1+\xi^{4})}{(1+\xi^{2})}$. For $\xi\ll 1$, $g(\xi)\sim\xi^{-2}>1$; while for $\xi\sim 1$ (i.e. $\xi=1-\varepsilon$, $\varepsilon\ll 1$), $g(\xi)\sim(1-\varepsilon)^{-2}>1$. Thus, we have that $g(\xi<1)>1$ and $g(\xi=1)=1$; or, in other words, $g(\xi=1)=1$ is the minimum value of $g(\xi)$ in the interval $\xi\in(0,1]$. This minimum gives the threshold value for the tension
$T_{0}^{c}\sim\frac{B}{W^{2}}.$ (S31)
@@
In terms of the longitudinal in-plane end-stretching imposed, $\epsilon_{0}$, we conclude that the end-stretched sheet buckles when the longitudinal in-plane displacement $\epsilon_{0}$ exceeds $\epsilon_{0}^{c}$, with
$\epsilon_{0}^{c}\sim\frac{B}{YW^{2}}.$ (S32)
Appendix F Effect of vanishing Poisson’s ratio
We examine the limit of vanishing Poisson’s ratio $\nu\to 0$ in the tension-induced transverse actuation of a sheet. In Fig. S12 (upper row) we plot the transverse component of the stress tensor (S10a) for a localized-loaded sheet, $W_{\mathrm{load}}\ll W$. Over a very wide range of Poisson’s ratios, $0.003\leq\nu\leq 0.3$, the results show that the compression zone — depicted in blue — persists when $\nu\to 0$. We emphasize that the same color scale is used in these plots and, hence, that the magnitude of this compression is also maintained as $\nu\to 0$. Conversely, when we plot this same transverse component for a full-width loaded sheet, $W_{\mathrm{load}}=W$, both the compressed area and the maximum compression shrink as $\nu\to 0$ (see Fig. S12 bottom row, for which we use Eq. (9b) in Ref. Xin and Davidovitch (2021a)).
This comparison hints that the loading geometry is key to the localized TUG-folding demonstrated in the main text and distinguishes it from the tensional wrinkling of Cerda and Mahadevan Cerda and Mahadevan (2003). This robustness to varying $\nu$ suggests a geometric origin to the compression. To gain insight into the mechanism that converts the longitudinal tension into a transverse compression for the different geometries, we consider a reductionist model of the sheet (see e.g. the kinematic analysis in the model by Celli et al. Celli et al. (2023)). Consisting of four rigid right-angled triangles whose acute-angled vertices are connected with a hinge (see Fig. S12, leftmost panels), the model inscribes a void rectangle that allows for ‘rigid’ deformations when subject external loads. In this construction, a uniaxial point-force generates a torque with respect to the hinges outside the loading axis, and hence the four triangles rotate about them (see Fig. S12, leftmost upper panel), generating transverse compression from uniaxial stretching. However, for a full-width loading the net torque vanishes, there is no rotation (Fig. S12, leftmost lower panel) and hence no transverse compression.
Translating this heuristic model into the rectangular elastic sheet, a point loading converts into transverse compression by ‘rotation’ of the bulk material. Thus, the compression persists despite vanishing Poisson’s ratio. By contrast, the full-width loading stretches the bulk material longitudinally, and the transverse compression is hence a Poisson effect that vanishes when $\nu\to 0$.
Extended Data Fig. S12: Transverse stress in the planar state for different Poisson’s ratios. Leftmost panels: Reductionist model of a longitudinally-loaded sheet consisting of four rigid right-angled triangles whose acute-angled vertices are connected with a hinge, thus forming a rectangular sheet — in absence of external loading — that inscribes a void rectangle. Upper row: Localized-loaded sheet. Lower row: Full-width loaded sheet. From left to right: Density plots of the transverse stress at increasing Poisson’s ratio ($\nu=0.003,0.01,0.03,0.1,0.3$ as per the labels) for a localized-loaded sheet (upper row) and a full-width loaded sheet (lower row). Results for the localized loading are by the expression (S10a) (the longitudinal coordinate, $y$, is shifted $y\to y+1$). For the full-width loading, the planar stress is derived from the Airy potential Eq. (9b) in Ref. Xin and Davidovitch (2021a) (after normalizing its coefficients by the corresponding Poisson’s ratio in each case). The same colorbar (only shown in the rightmost panels of each row) applies to all panels within a row. Contour lines are displayed at $\sigma_{xx}/T_{0}=(0,-0.05,-0.1,-0.15)$ (localized-loaded sheet) and at $\sigma_{xx}/T_{0}=(0,-0.0001,-0.0005,-0.001,-0.002,-0.003,-0.004,-0.005)$ (full-width loaded sheet). The aspect ratio is $W/L=1/4$ and we sum $n=32$ terms in the Airy potential’s series to derive the transverse stress in each panel.
Appendix G Wrinkling at larger strain
The analysis of the main text showed that a sheet subject to a localized ‘tug’ force buckles and folds at smaller strains than does a sheet clamped over its full-width does. However, as the strain increases the width of the doubly curved region, denoted $\omega$, shrinks (see Eq. (9) of the main text) and so must, at some strain, become comparable to the loading width $W_{\mathrm{load}}$. At this point we expect the situation to change since, within this doubly-curved region at least, the loading is no longer ‘point-like’. In particular, in this region, and for large strains, the problem becomes analogous to the CM problem Cerda and Mahadevan (2003) and wrinkles may be expected to form.
We demonstrate this behavior using FEM simulations of a rectangular sheet subject to increasing tensile strains with $W_{\mathrm{load}}/W=0.4$, see Fig. S13. Notably, higher modes in the out-of-plane displacement emerge when $W_{\mathrm{load}}\simeq\lambda_{\mathrm{CM}}$, where
$\lambda_{\mathrm{CM}}=2\sqrt{\pi}\left(\frac{BL^{2}}{Y\bar{\epsilon}}\right)^{1/4}$
is the wavelength in the CM problem, Eq. (5) of Ref. Cerda and Mahadevan (2003), and we have used $\bar{T}=Y\bar{\epsilon}$ to eliminate tension. A lower (asymmetric) mode would in principle be stable at smaller strains, i.e. when $W_{\mathrm{load}}\ <\lambda_{\mathrm{CM}}$, yet we do not observe this in the simulations because we consider perfectly symmetric sheets—a small imperfection is only introduced to break the up-down symmetry.
Extended Data Fig. S13: Wrinkling at larger strain. Cross-sectional profiles $\zeta(x,y=0)$, obtained from FEM simulations, of a sheet subject to different longitudinal strains exerted by pulling clamps of width $W_{\mathrm{load}}/W=0.4$. The profiles at different imposed strains are shown by solid curves, with color used to distinguish profiles at different strains. Here, $W=20,L/W=3,t/W=1\times 10^{-3}$ and the load width is emphasized by the gray shaded region. We note that the localized TUG folding studied here gives way to wrinkles similar to those in the CM problem when the strain $\bar{\epsilon}$ becomes large enough: we use Eq. (5) of Ref. Cerda and Mahadevan (2003) to determine the expected wrinkle wavelength $\lambda_{\rm CM}$ and then plot $x=\pm 0.5\times\lambda_{\mathrm{CM}}(\bar{\epsilon})$ (dashed curves). Note that the transition to wrinkling occurs when $W_{\mathrm{load}}\approx\lambda_{\rm CM}$.